### v1

# Bochner’s technique in Einstein’s non-symmetric geometry

Abstract
A. Einstein considered a manifold with a non-symmetric (0,2)-tensor $G$ and a linear connection $\nabla=(\Gamma^{k}_{ij})$ satisfying the condition $\dfrac{\partial G_{ij}}{\partial x^{k}}=\Gamma^{p}_{ik}G_{pj}+\Gamma^{p}_{kj}G_{ip}$. Guided by the construction of an almost Lie algebroid (on a vector bundle), we define the following concepts of Bochner’s technique for the Einstein’s non-symmetric geometry: the $\nabla^{f}$-connection, Bochner and Hodge $f$-Laplacians on tensors, the $f$-curvature operator and the Weitzenböck type curvature operator of Einstein’s connection. We prove Weitzenböck type decomposition formula and obtain vanishing results about the null space of the Bochner and Hodge $f$-Laplacians.
Keywords: Einstein’s non-symmetric geometry, curvature operator, almost Lie algebroid; Hodge and Bochner Laplacians, Weitzenböck curvature operator; singular distribution.
Mathematics Subject Classifications (2010) 53C15; 53C21
1 Introduction
Since 1923, A. Einstein worked on various versions of the Unified Field Theory (Non-symmetric Gravitational Theory – NGT), see [3]. This theory was intended to unite the gravitation theory, to which General Relativity (GR) is related, and the theory of electromagnetism. Introducing various versions of his NGT, A. Einstein used a complex basic tensor, with a symmetric real part and a skew-symmetric imaginary part. Beginning in 1950, A. Einstein used a real non-symmetric basic tensor $G=(G_{ij})$, whose symmetric part $g=(g_{ij})$ is associated with gravity, and the skew-symmetric part $F=(F_{ij})$ is associated with electromagnetism. The same is true for the symmetric part of the connection and torsion tensor, respectively. In his attempt to construct an unified field theory, (Einstein’s Gravitational Theory, briefly NGT), A. Einstein [3] considered a manifold $(M,G=g+F)$ with a linear connection $\nabla=(\Gamma^{k}_{ij})$ (not uniquely) determined by the condition $\dfrac{\partial G_{ij}}{\partial x^{k}}=\Gamma^{p}_{ik}G_{pj}+\Gamma^{p}_{kj}G_{ip}$ (called an Einstein’s connection), or, in coordinate-free form, see [4, 16]:
$XG(Y,Z)-G(\nabla_{Y}X,Z)-G(Y,\nabla_{X}Z)-G([X,Y],Z)=0\quad(X,Y,Z\in\mathfrak{X}_{M}).$ (1)
A manifold $(M,G=g+F)$ equipped with an Einstein’s connection $\nabla$ different from the Levi-Civita connection $\nabla^{g}$ is called an NGT space, or Einstein’s non-symmetric geometry.
The idea of a non-symmetric basic tensor was later revisited by J.W. Moffat [6], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [5] suggest potential improvements and challenges within NGT. M.I. Wanas et al. [15] studied path equations of a particle in Einstein’s non-symmetric geometry. Recent approaches to modified gravity often rely on differential geometry, incorporating torsion and non-metricity as natural extensions of GR framework. S. Ivanov and M. Zlatanović [4] investigated NGT using connections with skew-torsion and gave examples with the skew-symmetric part $F$ of $G$. V. Rovenski and M. Zlatanović [16] were the first to apply weak $f$-structures, in particular, weak almost contact structures, to NGT of skew-torsion with $F(X,Y)=g(X,fY)$. Weak metric structures (developed in recent works by V. Rovenski, see surveys [13, 14]) generalize the classical structures and are well suited for a study of $G=g+F$ in NGT with an arbitrary skew-symmetric tensor $F$.
The Bochner technique in differential geometry and mathematical physics is a powerful analytic method used to prove vanishing theorems by relating topology of a complete Riemannian manifold to its curvature. The Bochner technique works for skew-symmetric tensors lying in the kernel of the Hodge Laplacian $\Delta_{H}=d\,\delta+\delta\,d$: using maximum principles, one proves that such tensors are parallel, e.g., [8, 12]. Here $d$ is the exterior differential operator, and $\delta$ is its adjoint operator for the $L^{2}$ inner product. The elliptic differential operator $\Delta_{H}$ decomposes into two terms,
$\Delta_{H}=\nabla^{*}\nabla+\Re,$ (2)
one is the Bochner Laplacian $\nabla^{*}\nabla$, and the second term (depends linearly on the Riemannian curvature tensor), is called the Weitzenböck curvature operator on $(0,k)$-tensors $S$, e.g., [8], where
$\Re\,(S)(X_{1},\ldots,X_{k})=\sum\nolimits_{a=1}^{k}\sum\nolimits_{\,i=1}^{n}(R_{\,e_{i},X_{a}}\,S)(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}).$ (3)
Here $\{e_{i}\}$ is a local orthonormal frame on $(M,G=g+F)$ and $\nabla^{*}$ is the $L^{2}$-adjoint of the Levi-Civita connection $\nabla$. Note that $\Re$ reduces to $\rm Ric$} when evaluated on (0,1)-tensors, i.e., $k=1$. According to the well-known formula $(R_{Z,Y}\,S)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,i}S(X_{1},\ldots R_{Z,Y}X_{i},\ldots,X_{k})$ for the action of the curvature tensor $R$ on $(0,k)$-tensors, for $k\geq 2$ the formula from (3) has the form
$\Re(S)(X_{1},\ldots,X_{k})=$ $-2\sum\nolimits_{\,i,j,a;b<a}R(e_{i},X_{a},e_{j},X_{b})\cdot S(\underbrace{X_{1},\ldots,e_{j}}_{b},\underbrace{\ldots,e_{i}}_{a-b},\ldots,X_{k})$
$+\,\sum\nolimits_{\,i,a}{\rm Ric}(e_{i},X_{a})\cdot S(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}).$
The Weitzenböck formula (2) allows us to extend the Hodge Laplacian to arbitrary tensors.
In [10, 11], we expanded the Bochner’s technique to a Riemannian manifold endowed with a singular or regular distribution and Levi-Civita or a statistical type connection, and generalized some results in [7]. Singularities play a crucial role in mathematics and its applications, e.g., in theories of nonlinear control systems and GR. A singular distribution $\cal D$} on a manifold $M$ assigns to each point $x\in M$ a linear subspace ${\cal D}_{x}$ of the tangent space $T_{x}M$ in such a way that, for any $v\in{\cal D}_{x}$, there exists a smooth vector field $V$ defined in a neighborhood $U$ of $x$ and such that $V(x)=v$ and $V(y)\in{\cal D}_{y}$ for all $y\in U$. A priori, the dimension $\dim{\cal D}_{x}$ depends on $x\in M$. If $\dim{\cal D}_{x}=\operatorname{const}$, then $\cal D$} is a regular distribution. The distribution ${\cal D}={f}(TM)$ is singular in the sense of the definition given above. Indeed, for any vector $v\in{f}(T_{x}M)$ there exists a vector $\tilde{v}\in T_{x}M$ such that ${f}(\tilde{v})=v$. Take a smooth vector field $\tilde{V}$ in a neighborhood $U$ of $x$ such that $\tilde{V}(x)=\tilde{v}$. Then V=f(V~)V={f}($\tilde${V}) is a smooth vector field in $U$ and such that $V(x)=v$ and $V(y)\in{f}(T_{y}M)$ for all $y\in U$.
In this paper, we continue our study [10, 11] and develop the Bochner technique for NGT spaces: the $f$-connection, $f$-divergence, $f$-curvature tensor, Bochner and Hodge $f$-Laplacians and Weitzenböck $f$-curvature operator (where $f$, see (5), is conjugate to $F$). Our results hold both in the case where $f$ is non-degenerate and in the case where $\ker f$ defines a singular distribution (i.e, $\dim\ker f>0$ at some points), which applies to the weak metric $f$-structure, see [13, 14].
The work has the Introduction and two sections, the References include 16 items. In Section 2, we recall the basics of Einstein’s non-symmetric geometry and the almost Lie algebroid structure. In Section 3 (consisting of five subsections), we define for a manifold $(M,G=g+F)$ with an Einstein’s connection $\nabla$, the derivatives $\nabla^{f}$ and $d^{f}$ (Section 3.1), the $f$-divergence and their $L^{2}$ adjoint operators on tensors (Section 3.2), and Bochner and Hodge $f$-Laplacians on tensors and differential forms (Section 3.3). In Section 3.4, making some assumptions about $f$, we define the $f$-curvature operator $R^{f}$ of $\nabla^{f}$, and in Section 3.5, we define the Weitzenböck type curvature operator on tensors, prove Weitzenböck type decomposition formula and obtain vanishing results.
2 Preliminaries
Here we recall the basics of Einstein’s NGT and the almost Lie algebroid structure.
2.1 Einstein’s non-symmetric geometry
The basic (0,2)-tensor $G=(G_{ij})$ in a “generalized” Riemannian manifold $(M,G)$ is, in general, non-symmetric. It decomposes in two parts, $G=g+F$, the symmetric part $g=(g_{ij})$ (called Riemannian metric) and the skew-symmetric part $F=(F_{ij})$ (called fundamental 2-form), where
$g(X,Y)=\frac{1}{2}\big[G(X,Y)+G(Y,X)\big],\quad F(X,Y)=\frac{1}{2}\big[G(X,Y)-G(Y,X)\big].$ (4)
We assume that the symmetric part, $g$, is non-degenerate of arbitrary signature, and the skew-symmetric part, $F\neq 0$, has arbitrary (not necessarily constant) rank, e.g., is non-degenerate. Therefore, we obtain a (1,1)-tensor ${f}\neq 0$ (of not necessarily constant rank) determined by
$g(X,fY)=F(X,Y)\quad(X,Y\in\mathfrak{X}_{M}).$ (5)
Since $F$ is skew-symmetric, the tensor f{f} is also skew-symmetric: g(fX,Y)=g(X,fY)g({f}X,Y)=-g(X,{f}Y).
We consider linear connections $\nabla$ on $M$ with a torsion (1,2)-tensor $T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$. We denote the torsion (0,3)-tensor with respect to $g$ by the same letter,
$T(X,Y,Z):=g(T(X,Y),Z).$ (6)
A linear connection on a manifold $(M,G=g+F)$ is completely determined by the torsion tensor and the covariant derivative $\nabla g$ of the symmetric part $g$ of $G$, see [4].
The covariant derivative an $(p,k)$-tensor $S$, where $p=0,1$, as an $(p,k+1)$-tensor $\nabla S$:
$(\nabla S)(Y,X_{1},\ldots,X_{k})=\nabla_{Y}(S(X_{1},\ldots,X_{k}))-\sum\nolimits_{\,i=1}^{k}S(X_{1},\ldots,\nabla_{Y}X_{i},\ldots,X_{k}).$ (7)
Using (4), (5), (6) and (7), the condition (1) can be presented in the following form, see [4]:
$(\nabla G)(X,Y,Z)=-G(T(X,Y),Z)\ \Longleftrightarrow\ (\nabla(g+F))(X,Y,Z)=-T(X,Y,Z)-T(X,Y,{f}Z).$ (8)
Separating symmetric and skew-symmetric parts of (8) (with respect to $Y$ and $Z$), we express the covariant derivatives $\nabla g$ and $\nabla F$ in terms of torsion:
$(\nabla\,g)(X,Y,Z)$ $=-\frac{1}{2}\big[T(X,Y,Z)+T(X,Z,Y)+T(X,Y,{f}Z)+T(X,Z,{f}Y)\big],$
$(\nabla\,F)(Z,X,Y)$ $=\frac{1}{2}\big[T(X,Z,Y)-T(X,Y,Z)-T(X,Y,{f}Z)+T(X,Z,{f}Y)\big].$
The Einstein’s connection $\nabla$ is represented in [4] as
$g(\nabla_{X}Y,Z)$ $=g(\nabla^{g}_{X}Y,Z)+\frac{1}{2}\big[T(X,Y,Z)+T(X,Z,{f}Y)+T(Y,Z,{f}X)\big].$ (9)
The difference (1,2)-tensor $K$ of a linear connection $\nabla$ and the Levi-Civita connection $\nabla^{g}$ is
$K_{X}Y:=\nabla_{X}Y-\nabla^{g}_{X}Y.$ (10)
By (9), the difference tensor and torsion of an Einstein’s connection $\nabla$ are related as
$K(X,Y,Z)=\frac{1}{2}\,\big[T(X,Y,Z)+T(X,Z,{f}Y)+T(Y,Z,{f}X)\big],$ (11)
where $K(X,Y,Z):=g(K_{X}Y,Z)$. Therefore, for an Einstein’s connection $\nabla$ we have
$K_{X}Y-K_{Y}X=T(X,Y).$ (12)
Remark 1.
Linear connections with skew-torsion, T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), have become popular due to the relations with mathematical physics. In this case, the number of preserved supersymmetries is equal to the number of parallel spinors with respect to such a connection. S. Ivanov S. and M.Lj. Zlatanović, see [4], presented conditions for the existence and uniqueness of the Einstein’s connection with skew-torsion on a manifold $(M,G=g+F)$ and gave its explicit expression. In this case, the torsion is determined by $dF$: $T(X,Y,Z)=-\frac{1}{3}dF(X,Y,Z)$, and the difference tensor is given by the following formula:
$K_{X}Y=\frac{1}{2}\big[T({f}X,Y)-T(X,{f}Y)+T(X,Y)\big].$ (13)
Example 1.
A weak almost Hermitian manifold $(M,f,Q,g)$ is a Riemannian manifold $(M,g)$ of dimension $2m\geq 4$ endowed with non-singular endomorphisms: $f$ (skew-symmetric) and $Q$ (self-adjoint), and the 2-form $F$ such that the following conditions are valid, see [16]:
$f^{2}=-Q,\quad g(fX,fY)=g(QX,Y),\quad F(X,Y)=g(X,fY).$ (14)
Let $(M,f,Q,g)$ be a weak nearly Kähler manifold, i.e., (14) and $(\nabla^{g}_{X}f)X=0$ are true, considered as a generalized Riemannian manifold $(M,G=g+F)$. Suppose that $\nabla$ is an Einstein’s connection on $(M,f,Q,g)$ with skew-torsion. Then $\nabla$ satisfies the $f$-torsion condition: $T(fX,Y)=T(X,fY)$, and torsion of $\nabla$ is determined by $dF$ as $T(X,Y,Z){=}-\frac{1}{3}dF(X,Y,Z)$, see Example 3.4 in [16]. In this case, (13) reduces to $K_{X}Y=\frac{1}{2}\,T(X,Y)$, and our Einstein’s connection $\nabla$ is recovered by the equality $\nabla_{X}Y=\nabla^{g}_{X}Y+\frac{1}{2}\,T(X,Y)$.
Remark 2.
A linear connection $\nabla$} on a Riemannian manifold $(M,g)$ is said to be statistical if it is torsionless and tensor ${\nabla}g$ is symmetric in all its entries, equivalently, the cubic form $K(X,Y,Z)$ (associated with the difference tensor of $\nabla$) is symmetric. The statistical structure has applications in theory of probability and statistics as well as in information geometry. B. Opozda [7] developed Bochner’s technique for the statistical structure on a Riemannian manifold. In [11], we generalized Bochner’s technique to a Riemannian manifold equipped with a regular or singular distribution and a statistical type connection. Unfortunately, we cannot directly apply the statistical structure to Einstein’s non-symmetric geometry: if the cubic form $K(X,Y,Z)$ is symmetric, then by (11) and (12), $T=K=0$ is true; hence, $\nabla=\nabla^{g}$.
For an Einstein’s connection $\nabla$ on a manifold $(M,G=g+F)$, the tensor $K(X,Y,Z)$ cannot be totally symmetric, see Remark 2. Thus, it is interesting to study its symmetries separately.
The skew-symmetry condition of $K_{X}$ that is, K(X,Y,Z)=K(X,Z,Y)K(X,Y,Z)=-K(X,Z,Y), by (11), reduces to
$T(X,Y,Z)+T(X,Z,Y)+T(X,Z,fY)+T(X,Y,fZ)=0.$ (15)
Using (15), we find that $K_{X},K_{Y}]:TM\to TM$ is a skew-symmetric endomorphism:
$[K_{X},K_{Y}](Z,W)=-[K_{X},K_{Y}](W,Z),\quad X,Y,Z,W\in\mathfrak{X}_{M}.$ (16)
Define the vector fields
$E:=\sum\nolimits_{\,i}K_{e_{i}}e_{i},\quad E^{*}:=\sum\nolimits_{\,i}K^{*}_{e_{i}}e_{i},$ (17)
on $M$, where $K^{*}_{X}$ is the conjugate tensor to $K_{X}$: $g(K^{*}_{X}Y,Z)=g(Y,K_{X}Z)$, and $\{e_{i}\}$ is a local orthonormal frame. Note that (15) yields the equality $E^{*}=-E$.
2.2 The almost Lie algebroid structure
Here, we briefly recall the construction of an almost Lie algebroid, following [9, 10]. Lie algebroids (and Lie groupoids) constitute an active field of research in differential geometry. Roughly speaking, an (almost) Lie algebroid is a structure, where one replaces the tangent bundle $TM$ of a manifold $M$ with a new smooth vector bundle $\pi_{\cal E}\colon{\cal E}\to M$ of rank $k$, i.e., with a fibre $\mathbb{R}^{k}$, over $M$ with similar properties. Many geometrical notions, which involve $TM$, were generalized to the context of Lie algebroids.
Definition 1.
[…]
$Y,X]_{\rho}=-[X,Y]_{\rho},\quad[X,\psi Y]_{\rho}=\rho(X)(\psi)Y+\psi[X,Y]_{\rho},\quad\rho([X,Y]_{\rho})=[\rho(X),\rho(Y)]$ (18)
[…]
$\mathcal{J}_{\rho}(X,Y,Z)=\sum\nolimits_{\,\mathrm{cycl.}}[X,[Y,Z]_{\rho}\,]_{\rho},$
which prevents a bracket to satisfy the Jacobi identity, is called the Jacobiator of the bracket.
By (18), third formula, we get $\rho\,\mathcal{J}_{\rho}=0$. An almost Lie algebroid is a Lie algebroid provided that $\mathcal{J}_{\rho}=0$. Lie algebroids deal with integrable distributions (foliations). Almost Lie algebroids are closely related to singular distributions, i.e., given by $\rho({\cal E})$, see [10].
Note that axiom (18), third formula, is equivalent to vanishing of the following operator:
$\mathfrak{D}^{\rho}(X,Y)=[\rho X,\rho Y]-\rho([X,Y]_{\rho}).$ (19)
[…]
$d^{\rho}\,\omega(X_{0},\ldots,X_{k})=\sum\nolimits_{\,i=0}^{k}(-1)^{i}(\rho X_{i})(\omega(X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{k}))$
+ $+\mathrel{\mathop{\sum}\nolimits_{\,0\leq i<j\leq k}}(-1)^{i+j}\omega([X_{i},X_{j}]_{\rho},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}),$ (20)
where $X_{0},\ldots,X_{k}\in{\mathfrak{X}}_{\cal E}$ and $\omega\in\Lambda^{k}({\cal E})$ for $k\geq 0$. For $k=0$, we have $d^{\rho}f(X)=(\rho X)(f)$, where $X\in{\mathfrak{X}}_{\cal E}$ and $f\in C^{\infty}(M)=\Lambda^{0}({\cal E})$. Recall that a skew-symmetric bracket defines uniquely an exterior differential $d^{\rho}$ on $\Lambda(M)$, and it gives rise to
– an almost Lie algebroid if and only if $(d^{\rho})^{2}f=0$ for $f\in C^{\infty}(M)$;
– a Lie algebroid if and only if $(d^{\rho})^{2}f=0$ and $(d^{\rho})^{2}\,\omega=0$ for $f\in C^{\infty}(M)$ and $\omega\in\Lambda^{1}(M)$.
Definition 2.
A $\rho$-connection on $({\cal E},\rho)$ is a map $\nabla^{\rho}:\mathfrak{X}_{\cal E}\times\mathfrak{X}_{\cal E}\to\mathfrak{X}_{\cal E}$ satisfying Koszul conditions
$\nabla^{\rho}_{X}\,(fY+Z)=\rho(X)(f)Y+f\nabla^{\rho}_{X}\,Y+\nabla^{\rho}_{X}\,Z,\quad\nabla^{\rho}_{fX+Z}\,Y=f\nabla^{\rho}_{X}\,Y+\nabla^{\rho}_{Z}\,Y.$ (21)
[…]
$T^{\rho}(X,Y)=\nabla^{\rho}_{X}\,Y-\nabla^{\rho}_{Y}\,X-[X,Y]_{\rho},\quad R^{\rho}_{X,Y}=\nabla_{X}^{\rho}\nabla^{\rho}_{Y}-\nabla^{\rho}_{Y}\nabla^{\rho}_{X}-\nabla^{\rho}_{\,[X,Y]_{\rho}}.$
The following equality holds:
$\sum\nolimits_{\,\mathrm{circ.}}R^{\rho}_{\,X,Y}Z=\sum\nolimits_{\,\mathrm{circ.}}[(\nabla^{\rho}_{X}\,T^{\rho})(Y,Z)+T^{\rho}(T^{\rho}(X,Y),Z)]+\mathcal{J}_{P}(X,Y,Z).$
We will apply the construction of an almost Lie algebroid to a manifold $(M,G=g+F)$ with an Einstein’s connection, namely, in the rest of paper, we assume ${\cal E}=TM$ and $\rho=f$, see (5).
3 Main results
Here, motivated by the structure of an almost Lie algebroid, we develop Bochner’s technique for NGT spaces.
3.1 The modified covariant derivative and bracket
Definition 3.
Given an Einstein’s connection $\nabla$, on a manifold $(M,G=g+F)$, define the map $\nabla^{f}:\mathfrak{X}_{M}\times\mathfrak{X}_{M}\to\mathfrak{X}_{M}$:
$\nabla^{f}_{X}\,Y:=\nabla^{0}_{X}\,Y+K_{X}Y,$ (22)
called f{f}-connection. The $\nabla^{f}$ depends on the map $\nabla^{0}:\mathfrak{X}_{M}\times\mathfrak{X}_{M}\to\mathfrak{X}_{M}$ given by
$\nabla^{0}_{X}\,Y:=\nabla^{g}_{{f}X}\,Y,$
and the difference tensor $K$, see (10), but generally $\nabla^{f}$ and $\nabla^{0}$ are not linear connections on $M$. Set $\nabla^{f}_{X}\,\psi=({f}X)\psi=d\psi({f}X)$ for the f{f}-gradient of $\psi\in C^{1}(M)$. The f{f}-derivative of an $(p,k)$-tensor $S$, where $p=0,1$, is the following $(p,k+1)$-tensor $\nabla^{f}S$ :
$(\nabla^{f}S)(Y,X_{1},\ldots,X_{k}):=\nabla^{f}_{Y}(S(X_{1},\ldots,X_{k}))-\sum\nolimits_{\,i=1}^{k}S(X_{1},\ldots,\nabla^{f}_{Y}X_{i},\ldots,X_{k}).$ (23)
Such $\nabla^{f}$ (as well as $\nabla^{0}$) satisfies Koszul conditions (with any $\psi\in C^{\infty}(M)$), see (21),
$\nabla^{f}_{X}(\psi Y+Z)=({f}X)(\psi)Y+\psi\nabla^{f}_{X}\,Y+\nabla^{f}_{X}\,Z,\quad\nabla^{f}_{\psi X+Z}\,Y=\psi\nabla^{f}_{X}\,Y+\nabla^{f}_{Z}\,Y.$
We use the standard notation $\nabla^{f}_{Y}\,S=(\nabla^{f}S)(Y,\ldots)$. A tensor $S$ is called $\nabla^{f}$-parallel if $\nabla^{f}S=0$.
Using (23) and $\nabla^{g}\,g=0$, we obtain the following:
$(\nabla^{f}g)(X,Y,Z)$ $=-K(X,Y,Z)-K(X,Z,Y),$ (24)
$(\nabla^{f}F)(X,Y,Z)$ $=(\nabla^{g}\,F)({fX},Y,Z)-K(X,Y,Z)-K(X,Z,Y).$ (25)
For an Einstein’s connection $\nabla$, using (24), (25), and (11), we calculate
$(\nabla^{f}\,g)(X,Y,Z)$ $=-\frac{1}{2}\big[T(X,Z,Y)+T(X,Y,Z)+T(X,Y,fZ)+T(X,Z,{f}Y)\big],$
$(\nabla^{f}\,F)(X,Y,Z)$ $=(\nabla^{g}F)({fX},Y,Z)$
$\quad-\frac{1}{2}\big[T(X,Z,Y)+T(X,Y,Z)+T(X,Y,fZ)+T(X,Z,{f}Y)\big].$
The skew-symmetric f{f}-bracket $\cdot,\cdot]_{f}:\mathfrak{X}_{M}\times\mathfrak{X}_{M}\rightarrow\mathfrak{X}_{M}$ is defined by the formula
$X,Y]_{f}:=\nabla^{f}_{X}\,Y-\nabla^{f}_{Y}\,X.$ (26)
By this, the $f$-connection $\nabla^{f}$ is torsionless and satisfies conditions (18) with $\rho=f$.
Following (19) with $\rho=f$, we define the operator $\mathfrak{D}^{f}$ by the formula
$\mathfrak{D}^{f}(X,Y):=[{f}X,{f}Y]-{f}\,[X,Y]_{f}.$
Proposition 1.
For a manifold $(M,G=g+F)$ with an Einstein’s connection $\nabla$, the condition $\mathfrak{D}^{f}=0$ is equivalent to the following:
$(\nabla^{g}f)({f}X,Y)-(\nabla^{g}f)({f}Y,X)=f\,T(X,Y),\quad X,Y\in\mathcal{X}_{M}.$ (27)
Proof.
Put $X,Y]_{0}:=\nabla^{0}_{X}\,Y-\nabla^{0}_{Y}\,X$. Using (26), (22) and (12), we have
$[X,Y]_{f}$ $=\nabla^{g}_{fX}Y-\nabla^{g}_{fY}X+K_{X}Y-K_{Y}X=\nabla^{g}_{fX}Y-\nabla^{g}_{fY}X+T(X,Y)$
$=[X,Y]_{0}+T(X,Y).$ (28)
Using (3.1), we find
$\mathfrak{D}^{f}(X,Y)\hskip-0.28453pt$ $=$ $\hskip-0.28453pt\nabla^{g}_{fX}(fY)-f\nabla^{g}_{fX}Y-\nabla^{g}_{fY}(fX)+f\nabla^{g}_{fY}X-f\,T(X,Y)$
$=$ $\hskip-0.28453pt(\nabla^{g}f)({fX},Y)-(\nabla^{g}f)({fY},X)-f\,T(X,Y),$
and the required (27) follows. ∎
For an f{f}-connection $\nabla^{f}$ on $(M,G=g+F)$, its dual f{f}-connection $\breve{\nabla}^{f}$ is defined by the equality
${f}X(g(Y,Z))=g(\nabla^{f}_{X}Y,Z)+g(Y,\breve{\nabla}^{f}_{X}Z).$
In turn, the f{f}-connection $\nabla^{f}$ is dual to $\breve{\nabla}^{f}$. One may show that the following equality:
$\breve{\nabla}^{f}_{X}Y=\nabla^{0}_{X}Y-K^{*}_{X}Y$ (29)
is true in general. For a dual f{f}-connection $\breve{\nabla}^{f}$, we define the f{f}-bracket and $\breve{\mathfrak{D}}^{f}$ by
${[X,Y]}^{\smile}_{f}:=\breve{\nabla}^{f}_{X}\,Y-\breve{\nabla}^{f}_{Y}\,X,\quad\breve{\mathfrak{D}}^{f}(X,Y):=[{f}X,{f}Y]-{f}{[X,Y]}^{\smile}_{f}.$
If (15) holds, then using (29), we get $\breve{\nabla}^{f}_{X}Y=\nabla^{f}_{X}Y$; hence, the following is true:
${[X,Y]}^{\smile}_{f}=[X,Y]_{f},\quad\breve{\mathfrak{D}}^{f}=\mathfrak{D}^{f}.$ (30)
For any $k$-form $\omega$, set
$(K_{Y}\,\omega)(X_{1},X_{2},\ldots,X_{k}):=-\sum\nolimits_{\,i}\omega(X_{1},\ldots,K_{Y}X_{i},\ldots,X_{k}).$ (31)
The following lemma is similar to Lemmas 6.2 and 6.3 in [7].
Lemma 1.
Let $\nabla$ be an Einstein’s connection on a manifold $(M,G=g+F)$, $\omega\in\Lambda^{k}(M)$ and $\{e_{i}\}$ a local orthonormal frame. If condition (15) is true, then we have
$\sum\nolimits_{\,i}\iota_{\,e_{i}}((K_{e_{i}}+K^{*}_{e_{i}})\,\omega)=0,$ (32)
where $\iota$ stands for the interior product.
Proof.
We have
$\sum\nolimits_{\,i}((K_{e_{i}}+K^{*}_{e_{i}})\,\omega)(e_{i},X_{2},\ldots,X_{k})=$ $-\omega(\sum\nolimits_{\,i}(K_{e_{i}}+K^{*}_{e_{i}})e_{i},X_{2},\ldots,X_{k})$
$-\sum\nolimits_{\,i}\omega(e_{i},X_{2},\ldots,(K_{e_{i}}+K^{*}_{e_{i}})X_{p},\ldots,X_{k}).$
If $k=1$, the last term in the above does not appear. If $k>1$, we fix $2\leq p\leq k$, and calculate
$\sum\nolimits_{\,i}\omega(e_{i},X_{2},\ldots,(K_{e_{i}}+K^{*}_{e_{i}})X_{p},\ldots,X_{k})$
$=\sum\nolimits_{\,i<j}\big\{g((K_{e_{i}}+K^{*}_{e_{i}})X_{p},e_{j})-g((K_{e_{j}}+K^{*}_{e_{j}})X_{p}\big\}\omega(e_{i},X_{2},\ldots,e_{j},\ldots,X_{k}).$
Using (11) and the assumption (15) completes the proof of (32). ∎
3.2 The f{f}-divergence and Bochner f{f}-Laplacian
Define the f{f}-divergence of a vector field $X$ on a manifold $(M,G=g+F)$ by
$\operatorname{div}_{f}X=\operatorname{trace}(Y{\to}\,\nabla^{f}_{Y}\,X).$ (33)
The divergence of a $(1,k)$-tensor $S$ is defined by $\operatorname{div}_{g}S=\operatorname{trace}(Y\to\nabla^{g}_{Y}S)$, that is
$(\operatorname{div}_{g}S)(X_{1},\ldots,X_{k})=\sum\nolimits_{\,i}g((\nabla^{g}\,S)({e_{i}},X_{1},\ldots,X_{k}),e_{i}),$ (34)
using a local orthonormal frame $\{e_{i}\}$. The divergence of a vector field $X$ can also be defined by
$d(\iota_{X}\,d\operatorname{vol}_{g})=({\rm div}_{g}\,X)\,d\operatorname{vol}_{g}.$ (35)
The “musical” isomorphisms $\sharp:T^{*}M\to TM$ and $\flat:TM\to T^{*}M$ are used for rank one tensors, e.g. if $\omega\in\Lambda^{1}(M)$ and $X\in{\mathfrak{X}}_{M}$ then $\omega(X)=g(\omega^{\sharp},X)=g(\omega,X^{\flat})=X^{\flat}(\omega^{\sharp})$.
In order to generalize the Stokes Theorem for distributions, we formulate the following.
Lemma 2.
On a manifold $(M,G=g+F)$ with an Einstein’s connection $\nabla$, the condition
$\operatorname{div}_{g}{f}=(E^{*})^{\flat}$ (36)
is equivalent to the following equality:
$\operatorname{div}_{f}X=\operatorname{div}_{g}({f}X),\quad X\in{\mathcal{X}}_{M}.$ (37)
Proof.
Using $fe_{i}=\sum\nolimits_{\,j}g({f}e_{i},e_{j})e_{j}$ and (34) with $S=f$, we obtain
$\sum\nolimits_{\,i}g(\nabla^{g}_{{f}e_{i}}X,e_{i})=\sum\nolimits_{\,i,j}g({f}e_{i},e_{j})g(\nabla^{g}_{e_{j}}X,e_{i})=\sum\nolimits_{\,i,j}g(e_{i},-{f}e_{j})g(\nabla^{g}_{e_{j}}X,e_{i})$
$=\sum\nolimits_{\,j}g(\nabla^{g}_{e_{j}}X,-{f}e_{j})=\sum\nolimits_{\,j}g({f}\nabla^{g}_{e_{j}}X,e_{j})=\operatorname{div}_{g}({f}X)-(\operatorname{div}_{g}{f})(X).$
Using this and the definition (22), we have
$\operatorname{div}_{f}X=\sum\nolimits_{\,i}\,g(\nabla^{g}_{{f}e_{i}}X+K_{e_{i}}X,e_{i})=\operatorname{div}_{g}({f}X)-(\operatorname{div}_{g}{f})(X)+g(X,E^{*}).$
From this and (36), the required statement follows. ∎
Lemma 2 allows us to extend the Divergence Theorem to NGT spaces.
Theorem 1.
Let an Einstein’s connection $\nabla$ on a compact manifold $(M,G=g+F)$ with boundary satisfy (36). Then for any vector field $X\in\mathfrak{X}_{M}$ we have
$\int_{M}(\operatorname{div}_{f}X)\,d\operatorname{vol}_{g}=\int_{\partial M}g(X,\,{f}\nu)\,\omega_{\,\hat{g}}$, (38)
where $\nu$ is the unit inner normal to $\partial M$. In particular, on a manifold $(M,G=g+F)$ without boundary, for any $X\in\mathfrak{X}_{M}$ with compact support, (38) reduces to the equality $\int_{M}(\operatorname{div}_{f}X)d\operatorname{vol}_{g}{=}0$.
The following pointwise inner products and norms for $(0,k)$-tensors will be used:
$g(S_{1},\,S_{2})=\sum\nolimits_{\,i_{1},\ldots,\,i_{k}}S_{1}(e_{i_{1}},\ldots,\,e_{i_{k}})\,S_{2}(e_{i_{1}},\ldots,\,e_{i_{k}}),\quad\|S\|=\sqrt{g(S,\,S)},$
while, for $k$-forms, we set
$g(\omega_{1},\,\omega_{2})=\sum\nolimits_{\,i_{1}<\ldots\,<i_{k}}\omega_{1}(e_{i_{1}},\ldots,\,e_{i_{k}})\,\omega_{2}(e_{i_{1}},\ldots,\,e_{i_{k}}).$
For $L^{2}$-product of compactly supported tensors on a Riemannian manifold, we set
$(S_{1},\,S_{2})_{L^{2}}=\int_{M}g(S_{1},\,S_{2})\,d\operatorname{vol}_{g}.$
The $L^{2}$-adjoint operator $\nabla^{*{f}}$ for $\nabla^{f}$ maps $(p,k+1)$-tensors, where $p=0,1$, to $(p,k)$-tensors:
$(\nabla^{*{f}}S)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,i}(\nabla^{f}_{e_{i}}S)(e_{i},X_{1},\ldots,X_{k}),$ (39)
and similarly for $\breve{\nabla}^{*{f}}$.
Lemma 3.
The $\nabla^{*{f}}$ and $\breve{\nabla}^{*{f}}$ are related for any form $\omega\in\Lambda^{k}(M)$ by
$\breve{\nabla}^{*{f}}\omega=\nabla^{*{f}}\omega+\sum\nolimits_{\,i}\iota_{\,e_{i}}((K_{e_{i}}+K^{*}_{e_{i}})\,\omega);$ (40)
moreover, if (15) is true, then $\breve{\nabla}^{*{f}}\omega=\nabla^{*{f}}\omega$.
Proof.
Using (39) and (23), we relate $\nabla^{*{f}}$ and $\breve{\nabla}^{*{f}}$ with $\nabla^{*{0}}$ for any form $\omega\in\Lambda^{k}(M)$:
$\nabla^{*{f}}\,\omega=\nabla^{*{0}}\,\omega+\sum\nolimits_{\,i}\iota_{\,e_{i}}(K_{e_{i}}\,\omega),\quad\breve{\nabla}^{*{f}}\,\omega=\nabla^{*{0}}\,\omega-\sum\nolimits_{\,i}\iota_{\,e_{i}}(K^{*}_{e_{i}}\,\omega).$
Hence (40) is true. Applying (15) and (32) to (40), yields $\breve{\nabla}^{*{f}}\omega=\nabla^{*{f}}\omega$. ∎
The $\nabla^{*{f}}$ is related to the f{f}-divergence (33) by
$\operatorname{div}_{f}X=-\nabla^{*{f}}X^{\flat},\quad X\in{\cal X}_{M}.$ (41)
The next result shows when $\nabla^{*{f}}$ is $L^{2}$-adjoint to $\nabla^{f}$ on $k$-forms.
Proposition 2.
Let an Einstein’s connection $\nabla$ satisfy (15) and (36). Then for any compactly supported form $\omega_{1}\in\Lambda^{k}(M)$ and a form $\omega_{2}\in\Lambda^{k+1}(M)$, we have
$(\nabla^{*{f}}\omega_{2},\ \omega_{1})_{L^{2}}=(\omega_{2},\ \nabla^{f}\omega_{1})_{L^{2}}.$ (42)
Proof.
Define a compactly supported 1-form $\omega$ by
$\omega(Y)=g(\iota_{\,Y}\,\omega_{2},\,\omega_{1}),\quad Y\in{\mathcal{X}}_{M}.$
To simplify calculations, assume that $k=1$. Take an orthonormal frame $(e_{i})$ such that $\nabla_{Y}e_{i}=0$ for $Y\in T_{x}M$. To simplify calculations, assume that $s=t=1$, then at $x\in M$,
$-\nabla^{*{f}}\omega$ $=\sum\nolimits_{j}(\nabla^{f}_{e_{j}}\omega)(e_{j})=\sum\nolimits_{i,j}g({f}e_{j},e_{i})\,(\nabla_{e_{i}}\omega)(e_{j})=\sum\nolimits_{i,j,c}g({f}e_{j},e_{i})\,\nabla_{e_{i}}g(\omega_{2}(e_{j},e_{c}),\omega_{1}(e_{c})\,)$
$=\sum\nolimits_{i,j,c}g({f}e_{j},e_{i})\Big[g(\nabla_{e_{i}}\omega_{2}(e_{j},e_{c}),\omega_{1}(e_{c})\,)+g(\omega_{2}(e_{j},e_{c}),\nabla_{e_{i}}\omega_{1}(e_{c})\,)\Big]$
$=\sum\nolimits_{i,j,c}\Big[g(\,g({f}e_{j},e_{i})\nabla_{e_{i}}\omega_{2}(e_{j},e_{c}),\omega_{1}(e_{c}))+g(\omega_{2}(e_{j},e_{c}),g({f}e_{j},e_{i})\nabla_{e_{i}}\omega_{1}(e_{c})\,)\Big]$
$=\sum\nolimits_{j,c}\Big[g(\nabla^{f}_{e_{j}}\omega_{2}(e_{j},e_{c}),\omega_{1}(e_{c})\,)+g(\omega_{2}(e_{j},e_{c}),\nabla^{f}_{e_{j}}\omega_{1}(e_{c})\,)\Big]$
$=-g(\nabla^{*{f}}\omega_{2},\omega_{1})+g(\omega_{2},\nabla^{f}\omega_{1}).$
By the above equality, using (40), we obtain
$\nabla^{*{f}}\omega=-g(\breve{\nabla}^{*{f}}\omega_{2},\ \omega_{1})+g(\omega_{2},\ \nabla^{f}\omega_{1})-g(\iota_{E+E^{*}}\,\omega_{2},\,\omega_{1}).$ (43)
By (15), $E+E^{*}{=}0$ and $\breve{\nabla}^{*{f}}{=}\nabla^{*{f}}$ hold. Thus, by (41) and Theorem 1 with $X^{\flat}=\omega$, we get (42). ∎
The differential operator $\nabla^{*{f}}\,\nabla^{f}$ will be called the Bochner f{f}-Laplacian for an f{f}-connection.
Proposition 3.
Let an Einstein’s connection $\nabla$ on a closed manifold $(M,G=g+F)$ satisfy (15) and (36). If $g(\nabla^{*f}\,\nabla^{f}\omega,\,\omega)\leq 0$ holds for a form $\omega\in\Lambda^{k}(M)$, then $\nabla^{f}\omega=0$, i.e., $\omega$ is f{f}-parallel.
Proof.
We apply (42),
$0\geq(\nabla^{*{f}}\,\nabla^{f}\omega,\,\omega)_{L^{2}}=(\nabla^{f}\omega,\,\nabla^{f}\omega)_{L^{2}}\geq 0;$
hence, $\nabla^{f}\omega=0$. ∎
3.3 The Hodge f{f}-Laplacians
Using an Einstein’s connection $\nabla$, we define the exterior f{f}-derivative of a form $\omega\in\Lambda^{k}(M)$ by
$d^{f}\omega(X_{0},\ldots,X_{k})=\sum\nolimits_{\,i}(-1)^{i}(\nabla^{f}\omega)(X_{i},X_{0},\ldots,X_{i-1},X_{i+1},\ldots X_{k}),$ (44)
and similarly for $\breve{d}^{\,{f}}$. For $\psi\in C^{\infty}(M)$ and $X\in\mathfrak{X}_{M}$, we have $d^{f}\psi(X)=\nabla^{f}_{X}\psi$ and $\breve{d}^{\,{f}}\psi(X)=\breve{\nabla}^{f}_{X}\psi$.
Proposition 4.
The f{f}-derivative $d^{f}:{\mathrm{\Omega}}^{k}(M)\rightarrow{\mathrm{\Omega}}^{k+1}(M)$ is a 1-degree derivation, see (2.2), i.e.,
$d^{f}\omega(X_{0},\ldots,X_{k})=\sum\nolimits_{\,i=0}^{k}(-1)^{i}{f}X_{i}(\omega(X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{k}))$
+ $+\sum\nolimits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big([X_{i},X_{j}]_{f},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$ (45)
Similarly to (4), for any $\omega\in\Lambda^{k}(M)$, we get
${\breve{d}}^{f}\omega(X_{0},\ldots,X_{k})=\sum\nolimits_{\,i=0}^{k}(-1)^{i}{f}X_{i}(\omega(X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{k}))$
+ $+\sum\nolimits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big([X_{i},X_{j}]^{\smile}_{f},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$ (46)
Moreover, if condition (15) is valid, then ${\breve{d}}^{f}\omega=d^{f}\omega$.
Proof.
For any form $\omega\in\Lambda^{k}(M)$, using (44), we obtain
$d^{f}\omega(X_{0},\ldots,X_{k})=\sum\nolimits_{\,i=0}^{k}(-1)^{i}{f}X_{i}(\omega(X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{k}))$
$+\sum\nolimits_{\,i=0}^{k}(-1)^{i}\Big(\sum\nolimits_{\,j=0}^{i-1}\omega(X_{0},\ldots,\nabla^{f}_{X_{i}}X_{j},\ldots,X_{i-1},X_{i+1},\ldots,X_{k})$
$+\sum\nolimits_{\,j=i+1}^{k}\omega(X_{0},\ldots,X_{i-1},X_{i+1},\ldots,\nabla^{f}_{X_{i}}X_{j},\ldots,X_{k})\Big)$
$=\sum\nolimits_{\,i=0}^{k}(-1)^{i}{f}X_{i}(\omega(X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{k}))$
+ $+\sum\nolimits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big(\nabla^{f}_{X_{i}}X_{j}-\nabla^{f}_{X_{j}}X_{i},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$
Applying (26) to the above, we complete the proof of (4). The proof of (4) is similar. The last statement follows from the above and the first equality of (30). ∎
Define the f{f}-codifferential $\delta^{f}:\Lambda^{k}(M)\to\Lambda^{k-1}(M)$ by
$\delta^{\,f}:=\nabla^{*{f}}.$
Similarly, we set $\breve{\delta}^{\,f}:=\breve{\nabla}^{*{f}}$. Using the analog of (39), we obtain
$\breve{\delta}^{\,{f}}\omega(X_{2},\ldots,X_{k})=-\sum\nolimits_{\,i}(\breve{\nabla}^{f}_{e_{i}}\,\omega)(e_{i},X_{2},\ldots,X_{k}).$
Proposition 5.
On a closed manifold $(M,G=g+F)$ with an Einstein’s connection $\nabla$ satisfying conditions (15) and (36), for any forms $\omega_{1}\in\Lambda^{k}(M)$ and $\omega_{2}\in\Lambda^{k+1}(M)$, we have
$(\breve{\delta}^{\,{f}}\omega_{2},\,\omega_{1})_{L^{2}}=(\omega_{2},\,d^{f}\omega_{1})_{L^{2}}.$
Proof.
We derive as in the classical case:
$g(d^{f}\omega_{1},\,\omega_{2})$ $=\sum\nolimits_{\,u=0}^{k}(-1)^{i}\,\nabla^{f}_{\partial_{i_{u}}}\omega_{1}(\partial_{i_{0}},\ldots,\partial_{i_{u-1}},\partial_{i_{u+1}},\ldots,\partial_{i_{k}})g^{i_{0}j_{0}}\ldots g^{i_{k}j_{k}}\omega_{2}(\partial_{j_{0}},\ldots,\partial_{j_{k}})$
$=(k+1)\big(\nabla^{f}_{\partial_{i_{0}}}\omega_{1}(\partial_{i_{1}},\ldots,\partial_{i_{k}})\big)g^{i_{0}j_{0}}\ldots g^{i_{k}j_{k}}\omega_{2}(\partial_{j_{0}},\ldots,\partial_{j_{k}})=g(\nabla^{f}\omega_{1},\,\omega_{2}).$
It appears a $(k+1)$ factor, that finally is absorbed in the definition of $d^{f}$. Similarly, we get $g(\breve{d}^{f}\omega_{2},\,\omega_{1})=g(\breve{\nabla}^{*{f}}\omega_{2},\,\omega_{1})$. By the above, we have for the vector field $X=\omega^{\flat}$ (dual to the 1-form $\omega(Y)=g(\iota_{\,Y}\,\omega_{2},\,\omega_{1})$), an analog of (43):
$\operatorname{div}_{f}X=-g(\breve{d}^{f}\omega_{2},\,\omega_{1})+g(\omega_{2},\,d^{f}\omega_{1})+\sum\nolimits_{\,i}\iota_{\,e_{i}}((K_{e_{i}}+K^{*}_{e_{i}})\,\omega).$
Using this, (32) and Theorem 1, completes the proof. ∎
Definition 4.
Define the Hodge f{f}-Laplacians ${\mathrm{\Delta}}_{H}^{\,{f}}$ and $\breve{\mathrm{\Delta}}^{f}_{H}$ for differential forms by
${\mathrm{\Delta}}_{H}^{\,{f}}=d^{f}\breve{\delta}^{\,{f}}+\breve{\delta}^{\,{f}}d^{f},\quad\breve{\mathrm{\Delta}}^{f}_{H}=d^{f}\delta^{f}+\delta^{f}d^{f}.$
A differential form $\omega$ is called f{f}-harmonic if ${\mathrm{\Delta}}_{H}^{\,{f}}\,\omega=0$ and $\|\omega\|_{L^{2}}<\infty$ (similarly for $\breve{\mathrm{\Delta}}^{f}_{H}$).
The f{f}-harmonic forms have similar properties as in the classical case, e.g., [8]. Let (37) and (15) hold on a closed manifold $(M,G=g+F)$, then for $\omega\in\Lambda^{k}(M)$, by Proposition 5, we have
$({\mathrm{\Delta}}^{f}_{H}\,\omega,\,\omega)_{L^{2}}=(d^{f}\omega,\,d^{f}\omega)_{L^{2}}+(\breve{\delta}^{\,{f}}\omega,\,\breve{\delta}^{\,{f}}\omega)_{L^{2}};$
thus, such $\omega$ is f{f}-harmonic (and similarly for $\breve{\mathrm{\Delta}}^{f}_{H}$-harmonic case) if and only if $d^{f}\omega=0$ and $\breve{\delta}^{\,{f}}\omega=0$. If ${\mathrm{\Delta}}^{f}_{H}\,\omega=0$ and $\omega=d^{f}\,\theta$ hold for $\omega\in\Lambda^{k}(M)$, then $\breve{\delta}^{\,{f}}d^{f}\theta=\breve{\delta}^{\,{f}}\omega=0$; thus,
$(\omega,\,\omega)_{L^{2}}=(d^{f}\theta,\,d^{f}\theta)_{L^{2}}=(\theta,\,\breve{\delta}^{\,{f}}d^{f}\theta)_{L^{2}}=(\theta,\,\breve{\delta}^{\,{f}}\omega)_{L^{2}}=0.$
Moreover, if $\omega\in\Lambda^{k}(M)$ is f{f}-harmonic and $\omega=d^{f}\,\theta$ for some $\theta\in\Lambda^{k-1}(M)$, then $\omega=0$.
We also consider the Hodge type Laplacian related to $\nabla^{0}$, defined by $\Delta^{0}_{H}=\delta^{\,0}\,d^{\,0}+d^{\,0}\,\delta^{\,0}$, where for any differential form $\omega\in\Lambda^{k}(M)$,
$d^{\,0}\omega(X_{0},\ldots,X_{k})=\sum\nolimits_{\,i}(-1)^{i}(\nabla^{g}\omega)(fX_{i},X_{0},\ldots,X_{i-1},X_{i+1},\ldots X_{k}),$
$\delta^{\,0}\omega(X_{2},\ldots,X_{k})=-\sum\nolimits_{\,i}({\nabla}^{g}\omega)(fe_{i},e_{i},X_{2},\ldots,X_{k}).$
Lemma 4.
For an Einstein’s connection $\nabla$ and any form $\omega\in\Lambda^{k}(M)$ we have
[…]
Moreover, if condition (15) is valid, then
$\breve{\delta}^{f}\omega=\delta^{f}\omega,\quad\breve{\mathrm{\Delta}}^{f}_{H}\,\omega={\mathrm{\Delta}}^{f}_{H}\,\omega.$ (48)
Proof.
Using (23), (31) and (39), we obtain the first equality in (47):
$\delta^{f}\omega=-\sum\nolimits_{\,i}\nabla^{f}_{e_{i}}\iota_{e_{i}}\omega=\delta^{\,{0}}\omega-\sum\nolimits_{\,i}\iota_{e_{i}}(K_{e_{i}}\omega).$
Similarly, we obtain the second equality in (47). From this and (32) the equalities (48) follow. ∎
We will modify Stokes theorem on a complete open Riemannian manifold, see also Theorem 1.
Proposition 6.
Let a complete noncompact manifold $(M,G=g+F)$ be endowed with Einstein’s connection $\nabla$ and a vector field $X$ such that $\operatorname{div}_{f}X\geq 0$ or $\operatorname{div}_{f}X\leq 0$, and let conditions (36) and $\|{f}X\|_{g}\in\mathrm{L}^{1}(M,g)$ hold. Then $\operatorname{div}_{f}X\equiv 0$.
Proof.
Let $\omega$ be the (n1)(n-1)-form in $M$ given by $\omega=\iota_{{f}X}\,d\operatorname{vol}_{g}$, i.e., the contraction of the volume form $d\operatorname{vol}_{g}$ in the direction of fX{f}X. If $\{e_{1},\ldots,e_{n}\}$ is an orthonormal frame on an open set $U\subset M$ with coframe $\omega_{1},\ldots,\omega_{n}$}, then
$\iota_{{f}X}\,d\operatorname{vol}_{g}=\sum\nolimits^{n}_{i=1}(-1)^{i-1}g({f}X,e_{i})\,\omega_{1}\wedge\ldots\wedge\omega_{i-1}\wedge\omega_{i+1}\wedge\ldots\wedge\omega_{n}.$
Since the (n1)(n-1)-forms $\omega_{1}\wedge\ldots\wedge\omega_{i-1}\wedge\omega_{i+1}\wedge\ldots\wedge\omega_{n}$ are orthonormal in ${\mathrm{\Omega}}^{n-1}(M)$, we get $\|\omega\|_{g}^{2}=\sum\nolimits^{n}_{i=1}g({f}X,e_{i})^{2}=\|{f}X\|_{g}^{2}$. Thus, $\|\omega\|_{g}\in\mathrm{L}^{1}(M,g)$ and $d\omega(\operatorname{div}_{f}X)\,d\operatorname{vol}_{g}$, see (35) and (37). There exist domains $B_{i}$ on $M$ such that $M=\bigcup\nolimits_{\,i\geq 1}B_{i}$, $B_{i}\subset B_{i+1}$, and $\lim_{\,i\to\infty}\int\nolimits_{B_{i}}d\omega=0$. Then
[…]
But since $\operatorname{div}_{f}X\geq 0$ or $\operatorname{div}_{f}X\leq 0$ on $M$, it follows that $\operatorname{div}_{f}X=0$ on $M$. ∎
We call ${\mathrm{\Delta}}^{f}\psi=\operatorname{div}_{f}(\nabla^{f}\psi)$ the f{f}-Laplacian for a function $\psi\in C^{2}(M)$. By (22), we obtain
${\mathrm{\Delta}}^{f}\psi={{\mathrm{\Delta}}}^{0}\psi+({f}E)(\psi).$
The next theorem extends the well-known classical result on subharmonic functions.
Theorem 2.
Let a manifold $(M,G=g+F)$ be complete noncompact, and (36) hold for an Einstein’s connection $\nabla$. Assume that $\|{f}\nabla^{f}\psi\|_{g}$ and $\|\psi{f}\nabla^{f}\psi\|_{g}$ belong to $L^{1}(M,g)$ for some $\psi\in C^{2}(M)$ satisfying either ${\mathrm{\Delta}}^{f}\psi\geq 0$ or ${\mathrm{\Delta}}^{f}\psi\leq 0$. Then $\nabla^{f}\psi=0$; moreover, if either $f(TM)=TM$ (hence, $\dim M$ is even), or $f(TM)\neq TM$ and f(TM){f}(TM) is bracket-generating, then $\psi\!=\operatorname{const}$.
Proof.
Set $X=\nabla\,^{f}\psi$, then $\Delta^{f}\,\psi=\operatorname{div}_{f}X$. By Proposition 6 with $X=\nabla\,^{f}\psi$ and condition $\|{f}\nabla^{f}\psi\|\in L^{1}(M,g)$, we get $\Delta\,^{f}\,\psi\equiv 0$. Using
$\operatorname{div}_{f}(\psi\cdot Y)=\psi\cdot\operatorname{div}_{f}\,Y+g(\nabla^{f}\psi,\ Y)$
with $Y=\nabla\,^{f}\psi$, Proposition 6 with $X=\psi\nabla^{f}\psi$ and condition $\|\psi\,{f}\nabla\,^{f}\psi\|\in L^{1}(M,g)$, we get $(\nabla^{f}\psi,\,\nabla^{f}\psi)_{L^{2}}=0$, hence $\nabla^{f}\psi=0$. If $f(TM)=TM$, then $\psi\!=\operatorname{const}$. If $f(TM)\neq TM$, then $\nabla^{f}\psi=0$ means that $\psi\in C^{2}(M)$ is constant along the integral curves of $f(TM)$; moreover, if the singular distribution $f(TM)$ is bracket-generating then, by Chow’s theorem [2], $\psi={\rm const}$. ∎
3.4 The $f$-curvature operator
Definition 5.
Define the second f{f}-derivative of an $(p,k)$-tensor $S$ as the $(p,k+2)$-tensor
$(\nabla^{f})^{2}_{\,X,Y}\,S:=\nabla^{f}_{X}\,(\nabla^{f}_{Y}\,S)-\nabla^{f}\,\!\!_{\,\nabla^{f}_{X}Y}\,S,\quad X,Y\in\mathfrak{X}_{M}.$
Define the f{f}-curvature operator of $\nabla^{f}$ by standard formula, see Section 2.2,
$R^{f}_{X,Y}:=(\nabla^{f})^{2}_{\,X,Y}-(\nabla^{f})^{2}_{\,Y,X}=\nabla_{X}^{f}\,\nabla^{f}_{Y}-\nabla^{f}_{Y}\,\nabla^{f}_{X}-\nabla^{f}_{\,[X,Y]_{f}},\quad X,Y\in\mathfrak{X}_{M},$
and set
${R}^{f}(X,Y,Z,W)=g({R}^{f}_{X,Y}Z,\,W),\quad X,Y,Z,W\in\mathfrak{X}_{M}.$ (49)
The f{f}-Ricci curvature operator of $\nabla^{f}$ is defined by the standard way:
${\rm Ric}^{f}X=\sum\nolimits_{\,i}R^{f}_{\,X,e_{i}}\,e_{i},\quad{\rm Ric}^{f}(X,Y)=\sum\nolimits_{\,i}R^{f}(X,e_{i},e_{i},Y)=g({\rm Ric}^{f}X,Y).$ (50)
The action of $R^{f}$ on $(0,k)$-tensor fields is defined similarly to the action of the Riemannian curvature tensor $R^{g}$:
$(R^{f}_{X,Y}\,S)(X_{1},\ldots,X_{k})={\mathfrak{D}}^{f}(X,Y)(S(X_{1},\ldots,X_{k}))-\sum\nolimits_{\,i}S(X_{1},\ldots R^{f}_{X,Y}\,X_{i},\ldots,X_{k}).$ (51)
To simplify the calculations, in the rest of the paper we assume that the following Codazzi type condition is true:
$(\nabla^{g}_{{f}X}K)_{Y}Z=(\nabla^{g}_{{f}Y}K)_{X}Z,\quad X,Y,Z\in\mathfrak{X}_{M}.$ (52)
Here, $(\nabla^{g}_{{f}X}\,K)_{Y}\,Z=\nabla^{g}_{{f}X}(K_{Y}Z)-K_{\nabla^{g}_{{f}X}Y}\,Z-K_{Y}(\nabla^{g}_{{f}X}\,Z)$.
Proposition 7.
For an Einstein’s connection $\nabla$ with conditions (27) and (52), we have
  1. $(i)$
    ${R}^{f}_{X,Y}\,Z=R^{g}_{\,{f}X,{f}Y}\,Z+[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z$;
    $({R}^{f}_{X,Y}\,\omega)(Z)=-\omega({R}^{f}_{X,Y}\,Z-[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z)$;
    hence, $g({R}^{f}_{X,Y}\,Z,W)=-g({R}^{f}_{X,Y}\,W,Z)$,
  2. $(ii)$
    ${R}^{f}_{X,Y}\,\psi=0$ for any $\psi\in C^{2}(M)$;  if (15) holds, then ${R}^{f}_{X,Y}\,g=0$;
  3. $(iii)$
    for every $(1,k)$-tensor $S$ we have
    $({R}^{f}_{X,Y}\,S)(Z_{1},\ldots,Z_{k})=(R^{g}_{\,{f}X,{f}Y}\,S)(Z_{1},\ldots,Z_{k})$
    $+([K_{X},K_{Y}]$ ${-}K_{[X,Y]_{f}})S(Z_{1},\ldots,Z_{k})-\sum\nolimits_{\,i}S(Z_{1},\ldots([K_{X},K_{Y}]{-}K_{[X,Y]_{f}})Z_{i},\ldots,Z_{k}).$ (53)
  4. $(iv)$
    ${R}^{f}(X,Y,Z,W)=R^{g}({f}X,{f}Y,Z,W)+g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,W)$;
  5. $(v)$
    ${R}^{f}(X,Y,Z,W)=-{R}^{f}(Y,X,Z,W)$; if (15) holds, then ${R}^{f}(X,Y,Z,W)=-{R}^{f}(X,Y,W,Z)$.
Proof.
(i) Since ${f}{[X,Y]}_{f}=[{f}X,{f}Y]$ by the definition of ${\mathfrak{D}}^{f}$ and condition (27), we have
${R}^{f}_{X,Y}\,Z$ $=R^{g}_{\,{f}X,{f}Y}Z+(\nabla^{g}_{{f}X}\,K)_{Y}\,Z-(\nabla^{g}_{{f}Y}\,K)_{X}Z+[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,$
$({R}^{f}_{X,Y}\,\omega)(Z)$ $=-\omega(R^{g}_{\,{f}X,{f}Y}\,Z)+\omega((\nabla^{g}_{{f}X}\,K)_{Y}Z-(\nabla^{g}_{{f}Y}\,K)_{X}Z+[K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z).$
From this and (52) the first claim follows. Since (16) is true, that is, $K_{X},K_{Y}]:TM\to TM$ is skew-symmetric, ${R}^{f}_{X,Y}$ is also skew-symmetric: $g({R}^{f}_{X,Y}\,Z,W)=-g({R}^{f}_{X,Y}\,W,Z)$.
(ii) We calculate using our assumption (27), that is, ${\mathfrak{D}}^{f}=0$:
${R}^{f}_{X,Y}\,\psi={f}X({f}Y(\psi))-{f}Y({f}X(\psi))-({f}[X,Y]_{f})\psi={\mathfrak{D}}^{f}(X,Y)\psi=0.$
Next, using (i) we obtain
$g({R}^{f}_{X,Y}\,Z,W)=g(R^{g}_{\,{f}X,{f}Y}\,Z,W)+g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,\,W).$
Similarly, $g({R}^{f}_{X,Y}W,Z)=g(R^{g}_{\,{f}X,{f}Y}W,Z)+g([K_{X},K_{Y}]W-K_{[X,Y]_{f}}W,Z)$. By this and (51), we get
$({R}^{f}_{X,Y}\,g)(Z,W)=-g({R}^{f}_{X,Y}\,Z,W)-g(Z,{R}^{f}_{X,Y}\,W)$
$=(R^{g}_{\,{f}X,{f}Y}\,g)(Z,W)-g([K_{X},K_{Y}]Z,W)-g([K_{X},K_{Y}]W,Z)-g((K+K^{*})_{[X,Y]_{f}}Z,W)$
$=(R^{g}_{\,{f}X,{f}Y}\,g)(Z,W)-g((K+K^{*})_{[X,Y]_{f}}Z,W).$
Using $R^{g}_{\,{f}X,{f}Y}\,g=0$ and (15), we obtain ${R}^{f}_{X,Y}\,g=0$.
(iii) From the above and (51) the required result ($(iii)$) follows.
(iv) The equality follows from (49) and (i).
(v) Since ${R}^{f}_{X,Y}\,Z=-{R}^{f}_{Y,X}\,Z$, see (i), the first equality follows. For the second one, we use (15):
$g({R}^{f}_{X,Y}Z,\,Z)=0;$
thus, the claim follows from the equality $g({R}^{f}_{X,Y}(Z+W),\,Z+W)=0$. ∎
Similarly, we define the f{f}-curvature operator of the conjugate f{f}-connection $\breve{\nabla}^{f}$,
$\breve{R}^{f}_{X,Y}\,Z=\breve{\nabla}_{X}^{f}\,\breve{\nabla}^{f}_{Y}\,Z-\breve{\nabla}^{f}_{Y}\,\breve{\nabla}^{f}_{X}\,Z-\breve{\nabla}^{f}_{\,[X,Y]^{\smile}_{f}}Z,\quad X,Y,Z\in\mathfrak{X}_{M}.$
If we assume (15), i.e., $K^{*}_{X}=-K_{X}$, then using (29) we get $R^{f}_{X,Y}=\breve{R}^{f}_{X,Y}$.
Proposition 8.
For an Einstein’s connection $\nabla$ satisfying (15), (27) and (52), we have
${\rm Ric}^{f}(X,Y)={\rm Ric}^{0}(X,Y)-g(K_{X}E+\sum\nolimits_{\,i}K_{[X,e_{i}]_{f}}e_{i},\,Y)+g(K_{X},\,K_{Y}+T(Y,\,\cdot\,),$ (54)
where ${\rm Ric}^{{0}}(X,Y):=\sum\nolimits_{\,i}R^{g}({f}X,{f}e_{i},e_{i},Y)$ is the Ricci tensor of $\nabla^{{0}}$.
Proof.
Using (iv) of Proposition 7, (12) and (15), we have
${\rm Ric}^{f}(X,Y)$ $=\sum\nolimits_{\,i}{R}^{f}(X,e_{i},e_{i},Y)={\rm Ric}^{0}(X,Y)+\sum\nolimits_{\,i}g([K_{X},K_{e_{i}}]e_{i}-K_{[X,e_{i}]_{f}}e_{i},\ Y)$
$={\rm Ric}^{0}(X,Y)-g(K_{X}E,\,Y)+g(K_{X},\,K_{Y}+T(Y,\,\cdot\,)-\sum\nolimits_{\,i}K_{[X,e_{i}]_{f}}e_{i},\ Y),$
where $E$ is determined in (17). From the above the claim follows. ∎
Remark 3.
The skew-symmetric tensor f{f} induces a self-adjoint endomorphism $\cal P$} of $\Lambda^{2}(M)$:
${\cal P}(X\wedge Y)={f}X\wedge{f}Y,\quad{\cal P}^{*}(X\wedge Y)={f}^{*}X\wedge{f}^{*}Y\quad\Longrightarrow\quad{\cal P}^{*}={\cal P}.$
The curvature operator $R^{g}_{X,Y}$ can be seen as a self-adjoint linear operator ${\cal R}^{g}$ on the space $\Lambda^{2}(M)$ of bivectors, e.g., [1, 8]. Similarly, we consider $R^{f}_{X,Y}$ as a linear operator, or as a corresponding bilinear form on $\Lambda^{2}(M)$. For this, using our assumption (16) that the operator $K_{X},K_{Y}]:TM\to TM$ is skew-symmetric, define a linear operator $\cal K$} on $\Lambda^{2}(M)$ by
$g({\cal K}(X\wedge Y),Z\wedge W):=g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,\ W).$
Put ${\cal R}^{f}:={\cal R}^{g}\,{\cal P}+{\cal K}$ and $\breve{\cal R}^{f}:={\cal R}^{g}\,{\cal P}-{\cal K}$, i.e.,
${\cal R}^{f}(X\wedge Y)={\cal R}^{g}\,{\cal P}(X\wedge Y)+{\cal K}(X\wedge Y)={\cal R}^{0}({f}X\wedge{f}Y)+{\cal K}(X\wedge Y),$
${\cal R}^{f}(X\wedge Y,Z\wedge W)=g({\cal R}^{\,{f}}(X\wedge Y),Z\wedge W),$
$\breve{\cal R}^{f}(X\wedge Y)={\cal R}^{g}\,{\cal P}(X\wedge Y)-{\cal K}(X\wedge Y)={\cal R}^{g}({f}X\wedge{f}Y)-{\cal K}(X\wedge Y),$
$\breve{\cal R}^{f}(X\wedge Y,Z\wedge W)=g(\breve{\cal R}^{\,{f}}(X\wedge Y),Z\wedge W).$
The above ${\cal R}^{f}$ generalizes ${\cal R}^{0}:={\cal R}^{g}\,{\cal P}$ associated with $\nabla^{0}$, see [10]. Using known properties of ${\cal R}^{g}$ and property (iv) of $R^{f}$, see Proposition 7, we have
$g({\cal R}^{f}(X\wedge Y),Z\wedge W)=g({\cal R}^{g}({f}X\wedge{f}Y)+{\cal K}(X\wedge Y),\,Z\wedge W)=R^{f}(X,Y,W,Z),$
$g(\breve{\cal R}^{f}(X\wedge Y),Z\wedge W)=g({\cal R}^{g}({f}X\wedge{f}Y)-{\cal K}(X\wedge Y),\,Z\wedge W)=\breve{R}^{f}(X,Y,W,Z).$
Note that if ${\cal P}\,{\cal R}^{g}+{\cal K}^{*}\neq{\cal R}^{g}\,{\cal P}+{\cal K}$, then ${\cal R}^{f}$ on $\Lambda^{2}(M)$ is not self-adjoint:
$({\cal R}^{f})^{*}=({\cal R}^{g}\,{\cal P}+{\cal K})^{*}={\cal P}\,{\cal R}^{g}+{\cal K}^{*}\neq{\cal R}^{g}\,{\cal P}+{\cal K}={\cal R}^{f}.$
3.5 The Weitzenböck $f$-curvature operator
Here, we use an Einstein’s connection $\nabla$ to introduce the central concept of the Bochner technique: the Weitzenböck $f$-curvature operator on tensors. Namely, we generalize the Weitzenböck curvature operator (3), for the case of a manifold $(M,G=g+F)$ with an $f$-connection $\nabla^{f}$.
Definition 6.
Define the Weitzenböck $f$-curvature operator on $(0,k)$-tensors $S$ over $(M,G=g+F)$ by
$\Re^{f}(S)(X_{1},\ldots,X_{k})=\sum\nolimits_{\,a=1}^{k}\sum\nolimits_{\,i}({R}^{f}_{\,e_{i},X_{a}}\,S)(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}).$ (55)
The operators $\breve{\Re}^{\,{f}}$ and ${\Re}^{0}$ are defined similarly using $\breve{\nabla}^{f}$ and $\nabla^{0}$.
For a form $\omega\in\Lambda^{k}(M)$, the $\Re^{f}(\omega)$ is skew-symmetric, see (51) with ${\mathfrak{D}}^{f}=0$. For $k=1$, $\Re^{f}$ reduces to ${\rm Ric}^{f}$. For $k\geq 2$, using (51), the formula from (55) reads as
$\thinspace\Re^{f}(S)(X_{1},\ldots,X_{k})=-2\sum\nolimits_{\,i,j,a;b<a}R^{f}(e_{i},X_{a},e_{j},X_{b})\cdot S(\underbrace{X_{1},\ldots,e_{j}}_{b},\underbrace{\ldots,e_{i}}_{a-b},\ldots,X_{k})$
$+\,\sum\nolimits_{\,i,a}{\rm Ric}^{f}(e_{i},X_{a})\cdot S(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}).$ (56)
The following lemma represents $\Re^{f}$ using  $\Re^{\,{0}}$ and the difference tensor $K$.
Lemma 5.
Let an Einstein’s connection $\nabla$ on a manifold $(M,G=g+F)$ satisfy conditions (15), (27) and (52). Then
$\Re^{f}\omega=\Re^{\,{0}}\omega-\mathfrak{K}\,\omega,$ (57)
where the operator $\mathfrak{K}$ acts on forms $\omega\in\Lambda^{k}(M)$ by
$(\mathfrak{K}\,\omega)(X_{1},\ldots,X_{k})=2\sum\nolimits_{\,i,j,b<a}\big(g([K_{e_{i}},K_{X_{a}}]e_{j}-K_{[e_{i},X_{a}]_{f}}e_{j},\ X_{b})\big)\,\omega(\underset{b}{\underbrace{X_{1},\ldots,e_{j}}},\underset{a-b}{\underbrace{\ldots,e_{i}}}\ldots,X_{k})$
${-}\sum\nolimits_{\,a=1}^{k}\sum\nolimits_{\,i}\big(g(K_{e_{i}}E{+}\sum\nolimits_{\,j}\!K_{[e_{i},e_{j}]_{f}}e_{j},\,X_{a}){-}g(K_{e_{i}},\,K_{X_{a}}{+}T(X_{a},\,\cdot)\big)\,\omega(\underbrace{X_{1},\ldots,e_{i}}_{a},\ldots,X_{k}),$ (58)
for $k\geq 2$, and $(\mathfrak{K}\,\omega)(X){=}\sum\nolimits_{\,i}g(K_{e_{i}}E{+}\sum\nolimits_{\,j}\!K_{[e_{i},e_{j}]_{f}}e_{j},\,X){-}g(K_{e_{i}},\,K_{X}{+}T(X,\,\cdot))\omega(e_{i})$ for $k=1$.
Proof.
Using (iv) of Proposition 7 and (54), we find
${R}^{f}(e_{i},X_{a},e_{j},X_{b})$ $={R}^{0}(e_{i},X_{a},e_{j},X_{b})+g([K_{e_{i}},K_{X_{a}}]e_{j}-K_{[e_{i},X_{a}]_{f}}e_{j},X_{b}),$
${\rm Ric}^{f}(e_{i},X_{a})$ $={\rm Ric}^{\,{0}}(e_{i},X_{a})-g(K_{e_{i}}E{+}\sum\nolimits_{\,j}K_{[e_{i},e_{j}]_{f}}e_{j},\,X_{a}){+}g(K_{e_{i}},\,K_{X_{a}}{+}T(X_{a},\,\cdot\,).$ (59)
Substituting (3.5) in (3.5) and using linearity in the curvature and ($(iii)$) yields (57) and (5). ∎
The next theorem generalizes (2) and [10, Theorem 2] to the case of an Einstein’s connection.
Theorem 3.
Let an Einstein’s connection $\nabla$ satisfy conditions (15), (27) and (52). Then the following Weitzenböck type decomposition formula is valid for any form $\omega\in\Lambda^{k}(M)$:
$\Delta^{f}_{H}\,\omega=\breve{\nabla}^{*{f}}\,\nabla^{f}\,\omega+\Re^{f}(\omega).$ (60)
Proof.
Using the conditions, $E=E^{*}=0$, Lemma 3 and (48), we find
$d^{f}\,\breve{\delta}^{f}\omega(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,j}d^{f}\,\nabla_{e_{j}}^{\,{f}}\omega(e_{j},X_{1},\ldots,X_{k})$
$=\sum\nolimits_{\,j}\sum\nolimits_{a=0}^{k-1}(-1)^{a}\nabla_{X_{a+1}}^{\,{f}}\nabla_{e_{j}}^{\,{f}}\omega(e_{j},X_{1},\ldots X_{a},X_{a+2}\ldots,X_{k})$
$=-\sum\nolimits_{\,j}\sum\nolimits_{a=0}^{k-1}\nabla_{X_{a+1}}^{\,{f}}\nabla_{e_{j}}^{f}\omega\big(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}\big)$
$=-\sum\nolimits_{\,j,a}((\nabla^{f})^{2}_{X_{a+1},e_{j}}\,\omega)\big(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}\big)$
and
$\breve{\delta}^{f}\,d^{f}\omega(X_{1},\ldots,X_{k})=\nabla^{*{f}}(d^{f}\omega)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,j}\nabla_{e_{j}}^{f}(d^{f}\omega)(e_{j},X_{1},\ldots,X_{k})$
$=-\sum\nolimits_{\,j}\nabla^{f}_{e_{j}}\,\nabla^{f}_{e_{j}}\omega(X_{1},\ldots,X_{k})+\sum\nolimits_{\,j}\sum\nolimits_{a=0}^{k-1}(-1)^{a}\nabla^{f}_{e_{j}}\,\nabla^{f}_{X_{a+1}}\omega(e_{j},X_{1},\ldots,X_{a},X_{a+2},\ldots,X_{k})$
$=(\nabla^{*{f}}\,\nabla^{f}\omega)(X_{1},\ldots,X_{k})+\sum\nolimits_{\,j,a}((\nabla^{f})^{2}_{e_{j},X_{a+1}}\,\omega)(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}).$
From the above we have
$\Delta^{f}_{H}\,\omega=(\nabla^{*{f}}\,\nabla^{f}\omega)(X_{1},\ldots,X_{k})+\sum\nolimits_{\,j,a}\big((\nabla^{f})^{2}_{e_{j},X_{a+1}}\,\omega-(\nabla^{f})^{2}_{X_{a+1},e_{j}}\,\omega\big)(\underset{a+1}{\underbrace{X_{1},\ldots e_{j}}},\ldots,X_{k}).$
Thus, we have (60).
Example 2.
For vector fields and 1-forms, $\Re^{f}$ reduces to the kind of usual Ricci curvature, see (50) and (3.5). We have $\Re^{f}(\omega)(X)=\omega({\rm Ric}^{f}(X))$ for any $\omega\in\Lambda^{1}(M)$; thus, (60) reads as
$\Delta^{f}_{H}\,\omega=\breve{\nabla}^{*{f}}\,\nabla^{f}\omega+{\rm Ric}^{f}(\omega^{\sharp}).$
For every bivector $X\wedge Y\in\Lambda^{2}(M)$, we consider a map ${\cal R}^{f}(X\wedge Y):\mathfrak{X}_{M}\to\mathfrak{X}_{M}$, given by
$g({\cal R}^{f}(X\wedge Y)Z,W)\hskip-0.28453pt$ $=$ $\hskip-0.28453ptg({\cal R}^{f}(X\wedge Y),W\wedge Z)=R^{f}(X,Y,Z,W)$
$=$ $\hskip-0.28453ptR^{g}({f}X,{f}Y,Z,W)+g([K_{X},K_{Y}]Z,W),$
see Remark 3. Since bivectors are generators of the vector space $\Lambda^{2}(M)$, we obtain in this way a map ${\cal R}^{f}(\xi):\mathfrak{X}_{M}\to\mathfrak{X}_{M}$, where $\xi\in\Lambda^{2}(M)$ (similarly to algebraic curvature operator ${\cal R}^{g}(\xi)$).
Lemma 6.
For an Einstein’s connection $\nabla$ with conditions (15), (27) and (52), the map ${\cal R}^{f}(\xi)$, where $\xi\in\Lambda^{2}(M)$, is skew-symmetric: $g({\cal R}^{f}(\xi)W,\,Z)=-g({\cal R}^{f}(\xi)Z,\,W)$.
Proof.
It suffices to check the statement for the generators. Note that by condition (15), $K_{X}$ is skew-symmetric. Hence, using Proposition 7 and the conditions, we get
$g({\cal R}^{f}(X\wedge Y)Z,\ W)=R^{g}({f}X,{f}Y,Z,W)+g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,\ W)$
$=-R^{g}({f}X,{f}Y,W,Z)-g([K_{X},K_{Y}]W-K_{[X,Y]_{f}}W,\ Z)=-g({\cal R}^{f}(X\wedge Y)W,\ Z).$
Thus, the statement follows. ∎
The associated f{f}-curvature operator is given by
$g({\cal R}^{f}(X\wedge Y),\,Z\wedge W)=R^{g}({f}X,{f}Y,W,Z)-g([K_{X},K_{Y}]Z-K_{[X,Y]_{f}}Z,W).$
To simplify calculations, we assume that $\mathfrak{so}(TM)$ is endowed with metric induced from $\Lambda^{2}(M)$. If $L\in\mathfrak{so}(TM)$, then
$(L\,S)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,i}S(X_{1},\ldots,L(X_{i}),\ldots,X_{k}).$ (61)
Let $\{\xi_{a}\}$ be an orthonormal base of skew-symmetric transformations such that $(\xi_{a})_{x}\in\mathfrak{so}(T_{x}M)$ for $x$ in an open set $U\subset M$. By (61), for any $(0,k)$-tensor $S$,
$(\xi_{\alpha}S)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,i}S(X_{1},\ldots,\xi_{\alpha}(X_{i}),\ldots,X_{k});$
The ${\cal R}^{f}(X\wedge Y)$ on $\Lambda^{2}(M)$ can be decomposed using $\{\xi_{a}\}$.
Lemma 7.
For an Einstein’s connection $\nabla$ with conditions (15), (27) and (52), we have
${\cal R}^{f}(X\wedge Y)\hskip-0.28453pt$ $=$ $\hskip-0.28453pt-\sum\nolimits_{\,\alpha}\big(g({\cal P}\,{\cal R}^{g}(\xi_{\alpha})X,Y)+g({\cal K}(X\wedge Y),\xi_{\alpha})\big)\xi_{\alpha}$
$=$ $\hskip-0.28453pt-\sum\nolimits_{\,\alpha}\big(g({\cal R}^{g}(\xi_{\alpha}){f}X,{f}Y)+g({\cal K}(X\wedge Y),\xi_{\alpha})\big)\xi_{\alpha}.$
Proof.
Using $({\cal R}^{f})^{*}={\cal P}{\cal R}^{g}$ and Lemma 6, we have:
${\cal R}^{f}(X\wedge Y)\hskip-0.28453pt$ $=$ $\hskip-0.28453pt\sum\nolimits_{\,\alpha}g({\cal R}^{f}(X\wedge Y),\xi_{\alpha})\,\xi_{\alpha}$
$=$ $\hskip-0.28453pt\sum\nolimits_{\,\alpha}\big(g({\cal P}{\cal R}^{g}(\xi_{\alpha}),X\wedge Y)+g({\cal K}(X\wedge Y),\xi_{\alpha})\big)\xi_{\alpha}$
$=$ $\hskip-0.28453pt-\sum\nolimits_{\,\alpha}\big(g({\cal R}^{g}(\xi_{\alpha}){f}X,{f}Y)+g({\cal K}(X\wedge Y),\xi_{\alpha})\big)\xi_{\alpha},$
that completes the proof. ∎
Lemma 7 allows us to rewrite the Weitzenböck $f$-curvature operator (55).
Proposition 9.
For an Einstein’s connection $\nabla$ with conditions (15), (27) and (52), if $S$ is a $(0,k)$-tensor on $(M,G=g+F)$, then
$\Re^{f}(S)=-\sum\nolimits_{\,\alpha}{\cal R}^{f}(\xi_{a})(\xi_{a}S),\qquad(\Re^{f}(S))^{*}=\Re^{-{f}}(S).$ (62)
Proof.
We follow similar arguments as in the proof of [8, Lemma 9.3.3]:
$\quad\Re^{f}(S)(X_{1},\ldots,X_{k})=\sum\nolimits_{\,i,j}({\cal R}^{f}(e_{j}\wedge X_{i})S)(\underbrace{X_{1},\ldots,e_{j}}_{i},\ldots,X_{k})$
$=-\sum\nolimits_{\,i,j,\alpha}\big(g({\cal P}\,{\cal R}^{g}(\xi_{\alpha})e_{j},X_{i})+g({\cal K}(e_{j}\wedge X_{i}),\xi_{\alpha})\big)(\xi_{\alpha}S)(X_{1},\ldots,e_{j},\ldots,X_{k})$
$=-\sum\nolimits_{\,i,j,\alpha}(\xi_{\alpha}S)(X_{1},\ldots,\big(g({\cal P}{\cal R}^{g}(\xi_{\alpha})e_{j},X_{i})e_{j}+g({\cal K}(e_{j}\wedge X_{i}),\xi_{\alpha})\big),\ldots,X_{k})$
$=-\sum\nolimits_{\,i,j,\alpha}(\xi_{\alpha}S)(X_{1},\ldots,g(e_{j},{\cal R}^{f}(\xi_{\alpha})X_{i})e_{j},\ldots,X_{k})$
$=-\sum\nolimits_{\,i,\alpha}(\xi_{\alpha}S)(X_{1},\ldots,{\cal R}^{f}(\xi_{\alpha})X_{i},\ldots,X_{k})=-\sum\nolimits_{\,\alpha}({\cal R}^{f}(\xi_{\alpha})(\xi_{\alpha}S))(X_{1},\ldots,X_{k}).$`
Thus, (62)1 follows. Since ${\cal R}^{g}:\Lambda^{2}(M)\to\Lambda^{2}(M)$ is self-adjoint, there is a local orthonormal base $\{\xi_{a}\}$ of $\Lambda^{2}(M)$ such that ${\cal R}^{g}(\xi_{a})=\lambda_{a}\,\xi_{a}$. Using this base, for any $(0,k)$-tensors $S_{1}$ and $S_{2}$, we get
$g(\Re^{f}(S_{2}),\,S_{1})\hskip-0.28453pt$ $=$ $\hskip-0.28453pt-\sum\nolimits_{\,\alpha}g({\cal R}^{f}(\xi_{\alpha})(\xi_{\alpha}S_{2}),S_{1})=-\sum\nolimits_{\,\alpha}g(\xi_{\alpha}S_{2},\,({\cal R}^{f})^{*}(\xi_{\alpha})S_{1})$ (63)
$=$ $\hskip-0.28453pt\sum\nolimits_{\,\alpha}g(\xi_{\alpha}S_{2},\,({\cal P}{\cal R}^{g}+{\cal K})(\xi_{\alpha})(S_{1}))$
$=$ $\hskip-0.28453pt\sum\nolimits_{\,\alpha}\lambda_{\alpha}g({\cal P}(\xi_{\alpha}S_{2}),\,\xi_{\alpha}S_{1})+\sum\nolimits_{\,\alpha}g({\cal K}^{*}(\xi_{\alpha}S_{2}),\,\xi_{\alpha}S_{1}),$
and, similarly,
$g(S_{2},\,\Re^{-{f}}(S_{1}))=\sum\nolimits_{\,\alpha}\lambda_{\alpha}g(\,\xi_{\alpha}S_{2},\,{\cal P}(\xi_{\alpha}S_{1}))+\sum\nolimits_{\,\alpha}g(\xi_{\alpha}S_{2},\,{\cal K}(\xi_{\alpha}S_{1}))$
$=\sum\nolimits_{\,\alpha}\lambda_{\alpha}g({\cal P}(\xi_{\alpha}S_{2}),\,\xi_{\alpha}S_{1})+\sum\nolimits_{\,\alpha}g({\cal K}^{*}(\xi_{\alpha}S_{2}),\,\xi_{\alpha}S_{1}).$
Thus, (62)2 follows. ∎
The following result generalizes [8, Corollary 9.3.4].
Proposition 10.
Let $\nabla$ be an Einstein’s connection on a manifold $(M,G=g+F)$ satisfying conditions (15), (27) and (52).
a) If $g({\cal R}^{f}(S),S)\geq 0$ for any $(0,k)$-tensor $S$, then $g(\Re^{f}(S),S)\geq 0$.
b) If $g({\cal R}^{f}(S),S)\geq-\varepsilon\,\|S\|^{2}$ for any $(0,k)$-tensor $S$, where $\varepsilon>0$, then $g({\Re}^{\,{f}}(S),S)\geq-\varepsilon\,C\,\|S\|^{2}$,
where a constant $C$ depends only on the type of $S$.
Proof.
Using (63) and a local orthonormal base $\{\xi_{\alpha}\}$ of $\Lambda^{2}(M)$ such that ${\cal R}^{g}(\xi_{\alpha})=\lambda_{\alpha}\xi_{\alpha}$, we get
$g(\Re^{f}(S),S)\hskip-0.28453pt$ $=$ $\hskip-0.28453pt\sum\nolimits_{\,\alpha}\lambda_{\alpha}g({\cal P}(\xi_{\alpha}S),\,\xi_{\alpha}S)+\sum\nolimits_{\,\alpha}g({\cal K}(\xi_{\alpha}S),\,\xi_{\alpha}S)$
$=$ $\hskip-0.28453pt\sum\nolimits_{\,\alpha}g({\cal P}(\xi_{\alpha}S),\,{\cal R}^{g}(\xi_{\alpha}S))+\sum\nolimits_{\,\alpha}g({\cal K}(\xi_{\alpha}S),\,\xi_{\alpha}S)$
$=$ $\hskip-0.28453pt\sum\nolimits_{\,\alpha}g({\cal R}^{f}(\xi_{\alpha}S),\,\xi_{\alpha}S).$
By conditions, $g({\cal R}^{f}(\xi_{\alpha}S),\,\xi_{\alpha}S)\geq 0$ for all $\alpha$, thus, $g(\Re^{f}(S),S)\geq 0$, and the first claim follows. There is a constant $C>0$ depending only on the type of the tensor and $\dim M$ such that $C\|S\|^{2}\geq\sum\nolimits_{\,\alpha}\|\xi_{\alpha}S\|^{2}$, see [8, Corollary 9.3.4]. By conditions, $g({\cal R}^{f}(\xi_{\alpha}S),\,\xi_{\alpha}S)\geq-\varepsilon\,\|\xi_{\alpha}S\|^{2}$ for all $\alpha$. The above yields $g({\cal R}^{f}(\xi_{\alpha}S),\,\xi_{\alpha}S)\geq-\varepsilon\,C\,\|S\|^{2}$ – thus, the second claim. ∎
Theorem 4.
Let conditions (15), (27), (36) and (52) hold for an Einstein’s connection $\nabla$ on a closed manifold $(M,G=g+F)$. If $g({\cal R}^{f}(\omega),\omega)\geq 0$ for an f{f}-harmonic form $\omega\in\Lambda^{k}(M)$, then $\nabla^{f}\omega=0$.
Proof.
By conditions and Proposition 10(a), $g(\Re^{f}(\omega),\omega)\geq 0$ is true. By (60), since $\Delta^{f}_{H}\,\omega=0$, we get $g(\breve{\nabla}^{*{f}}\,\nabla^{f}\omega,\omega)\leq 0$. By Proposition 3, where (36) is used, we have $\nabla^{f}\omega=0$. ∎
Conclusion
The main contribution of this paper (four theorems and ten propositions) is the development of Bochner’s technique for a manifold $(M,G=g+F)$ with an Einstein’s connection. To generalize some geometrical analysis tools for an Einstein’s connection, we assume some additional conditions, for example, (15), (27), (36) and (52), for $F$ and the difference tensor $K$. We define the $f$-curvature operator and the Weitzenböck $f$-curvature operator for NGT spaces, and prove vanishing theorems on the null space of the Bochner and Hodge $f$-Laplacians. We delegate the following for further study: considering NGT spaces associated with special classes of weak metric $f$-manifolds, see [13, 14, 16], and find more applications in NGT and mathematical physics.
Acknowledgments. The author expresses his sincere gratitude to Prof. Milan Zlatanović for numerous fruitful discussions of NGT and its applications. His comments significantly improved both the accuracy and clarity of the manuscript.